OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)? -x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0

OpenStudy (anonymous):

@mathstudent55 ?

OpenStudy (anonymous):

y=x+7 its slope m=1 (y=mx+c) slope of line perpendicular to above line=-1/1=-1 (-1/m) eq. of any line with slope -1 is y=-1x+c (y=mx+c) because it passes through (-1,-1) so -1=-1*-1+c, c=-2 ?

OpenStudy (jdoe0001):

if you were to solve " -x + y = 7 " for "y", what would it look like?